OpenStudy (anonymous):

Simplify the expression. (2xy/y^-2)^-2

OpenStudy (ashleyisakitty):

http://www.webmath.com/_answer.php

OpenStudy (anonymous):

That didn't lead to anywhere?

OpenStudy (anonymous):

\[\left( \frac{ 2xy }{ y ^{-2} } \right)^{-2} \rightarrow \left( \frac{ 2xyy ^{2} }{ 1 } \right)^{-2} \rightarrow \left( 2xy^3 \right)^{-2}\rightarrow 2^{-2}x ^{-2}y ^{-6}\rightarrow \frac{ 1 }{ 4x ^{2}y ^{6} }\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (j-cross18):

Are you still here???